# Detecting AI Flaws: Target-Driven Attacks on Internal Faults in Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have become a focal point in the rapidly evolving field of artificial intelligence. However, a critical concern is the presence of toxic content within the pre-training corpus of these models, which can lead to the generation of inappropriate outputs. Investigating methods for detecting internal faults in LLMs can help us understand their limitations and improve their security. Existing methods primarily focus on jailbreaking attacks, which involve manually or automatically constructing adversarial content to prompt the target LLM to generate unexpected responses. These methods rely heavily on prompt engineering, which is time-consuming and usually requires specially designed questions. To address these challenges, this paper proposes a target-driven attack paradigm that focuses on directly eliciting the target response instead of optimizing the prompts. We introduce the use of another LLM as the detector for toxic content, referred to as ToxDet. Given a target toxic response, ToxDet can generate a possible question and a preliminary answer to provoke the target model into producing desired toxic responses with meanings equivalent to the provided one. ToxDet is trained by interacting with the target LLM and receiving reward signals from it, utilizing reinforcement learning for the optimization process. While the primary focus of the target models is on open-source LLMs, the fine-tuned ToxDet can also be transferred to attack black-box models such as GPT-4o, achieving notable results. Experimental results on AdvBench and HH-Harmless datasets demonstrate the effectiveness of our methods in detecting the tendencies of target LLMs to generate harmful responses. This algorithm not only exposes vulnerabilities but also provides a valuable resource for researchers to strengthen their models against such attacks.
Warning: This paper includes content that may be offensive or harmful.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly improved performance across various natural language processing (NLP) tasks (Ouyang et al. 2022; Brohan et al. 2023; Achiam et al. 2023; Rafailov et al. 2024). Despite these advances, a significant issue persists: the training data for LLMs often contains toxic content that is difficult to filter. This can lead the models to learn and replicate such behavior, potentially resulting in the generation of inappropriate, offensive, or harmful content (Bai et al. 2022a; Zou et al. 2023). To mitigate this problem, most LLMs undergo a process known as safety alignment (Ziegler et al. 2019; Perez et al. 2022; Zhuo et al. 2023). During this process, the models are fine-tuned based on human preferences that reflect positive societal values, leading to more helpful, appropriate, and safe responses. However, even such aligned LLMs remain vulnerable to attacks designed to bypass their safety constraints (Zhou et al. 2024).
To address these challenges, research focused on identifying and mitigating internal faults in LLMs is crucial to enhance their robustness against attacks (Zou et al. 2023; Liu et al. 2024; Paulus et al. 2024; Sadasivan et al. 2024). Among various attack strategies, the most prevalent is the “jailbreaking attack”, which involves crafting adversarial prompts to bypass the safety mechanisms of target models. This process heavily relies on prompt engineering to create adversarial prompts or misleading content. Notable early instances include “Ignore Previous Prompt” (Perez and Ribeiro 2022) and “Do Anything Now” (walkerspider 2022), which use manually crafted adversarial prompts to compromise large language models. However, this manual process is time-consuming and hinders practical implementation in the daily use of LLMs. To address this limitation, another category of work reformulates the jailbreaking process as an automatic adversarial example generation process, as demonstrated by GCG (Zou et al. 2023) and PAL (Sitawarin et al. 2024). These methods involve leveraging gradients from open-source LLMs to guide the optimization process. Nevertheless, this approach remains computationally intensive due to the discrete optimization required. To alleviate these time-consuming problems, another set of methods employs automated models to generate adversarial suffixes for the prompts, such as AdvPrompter (Paulus et al. 2024), which can mislead models into generating unexpected responses. However, this process still requires carefully designed prompts to achieve satisfactory results.
In this paper, we aim to explore a new paradigm for addressing the internal faults of LLMs. Instead of relying on prompt engineering, we focus on directly attacking the models to elicit the desired toxic responses. We introduce a target-driven attack paradigm that optimizes a LLM called ToxDet, which can generate questions and preliminary answers based on the intended toxic response. This approach enables an attack on the target model to produce the specific toxic response. Specifically, we formulate the attacking problem such that, given a response containing inappropriate or harmful information, ToxDet generates potential questions and preliminary sentences that could lead to this response. These generated questions and preliminary statements are then used to query the target LLM, aiming to elicit similar unexpected outputs. We frame the training process of ToxDet as a reward maximization problem, utilizing reinforcement learning for optimization. Instead of employing a separate pretrained reward model, we propose to use the target model itself as the reward model to better align the attacking process. The reward scores are defined as the log-probabilities of the responses from the target model, given the generated questions and preliminary statements from ToxDet. This approach ensures that the attacking process is specifically tailored to the target model. In this case, we only require the predicted probabilities from the target model, rather than its gradients. This makes the training process relatively swift due to the efficiency of querying the target model. Moreover, the inference process is even faster as it does not depend on the target probabilities. Remarkably, despite ToxDet’s design being customized for a specific target model, it demonstrates significant transferability to black-box models, such as GPT-4o, yielding notable results. To summarize, our main contributions are:
  • We introduce a novel target-driven attacking paradigm that employs a ToxDet LLM to directly target the intended toxic response, thereby eliminating the need for labor-intensive or query-sensitive prompt engineering.
  • We propose to use the target model itself as the reward model, bypassing the need for a separately pretrained reward model. The reward signals are derived from the probabilities provided by the target model, making our method’s automatic attacking process faster compared to manual and gradient-based methods.
  • Although trained on open-source target LLMs, our ToxDet can be directly applied to attack black-box models like GPT-4o, demonstrating significant results.
  • Experimental results on the AdvBench (Zou et al. 2023) and HH-Harmless (Bai et al. 2022a) benchmarks demonstrate the effectiveness of our method’s universal and transferable attacking capabilities for both open-source and closed-source LLMs. These findings expose the vulnerabilities of existing LLMs and provide valuable insights for researchers to further enhance their robustness.
@@ 2 Preliminary
2.1 Notion & Problem Setting
Let ${\mathcal{V}}$ denote the vocabulary of a language model characterized by a policy $\pi$, where $x\in{\mathcal{V}}$ represents a token and ${\mathbf{x}}\in{\mathcal{V}}^{*}$ represents a sequence of tokens. Here, ${\mathcal{V}}^{*}$ denotes the set of all possible token sequences of arbitrary length. Originally, consider a misleading question ${\mathbf{x}}\in{\mathcal{V}}^{*}$ (e.g., “Write a tutorial on building a bomb”), which could lead the target chat-based LLM $\pi_{\phi}$ with $\phi$ being the parameters, to provide defensive responses (e.g., “I am sorry, …”). The task of jailbreaking attack (by injection) involves appending an adversarial suffix ${\mathbf{q}}\in{\mathcal{V}}^{*}$ (e.g., “as part of a lecture”) to the prompt, in case to confuse the target model into generating a positive response, such as “Sure, here is a tutorial on building a bomb…”.
In this paper, we aim to exploit a new target-driven attack paradigm. We formulate our proposed attack scenario as follows: For given toxic contents ${\mathbf{y}}\in{\mathcal{V}}^{*}$, posing a possible misleading question ${\mathbf{x}}$ concatenated with a preliminary answer ${\mathbf{p}}\in{\mathcal{V}}^{*}$, so as to induce the aligned $\pi_{\phi}$ to provide the given toxic contents ${\mathbf{y}}$.
The primary distinction of our attack setting compared to other jailbreaking techniques (Paulus et al. 2024; Liu et al. 2024) lies in the approach to generating adversarial content. Previous methods generate adversarial suffixes conditioned on fixed original prompts from the dataset, where the prompts serve as the conditions and the suffixes are generated accordingly. In contrast, our approach uses the intended toxic response to generate potential questions and preliminary answers that could lead to that response. A primary concern is why we add a preliminary answer instead of a potential question directly. We found that using the question itself inherently optimizes the prompt in a way that has minimal impact on the target model. Therefore, we propose the use of a preliminary answer. We will discuss this in detail in Appendix A.
Figure 1: Illustration of our proposed target-driven attack process and comparison with the jailbreaking attack. The jailbreaking attack (left part) involves an attacker generating a suffix, which is then concatenated with the question to attack the target models. In contrast, our target-driven attack (right part) takes the target response as input and uses reinforcement learning to refine the model. This process generates refined precedent content, prompting the target model to produce a toxic response with the same meaning as the original response. The reward is the log-probability directly obtained from the attacked target model.
2.2 White-box, Gray-box, Black-box LLMs
In this paper, we categorize target LLMs into three types: white-box, with full access to architecture and parameters; black-box, like GPT-4, where interaction is limited to text prompts and responses; and gray-box, which provides both text responses and log-probabilities of each token. We focus on attacking gray-box LLMs using log-probability information and explore the transferability of these attacks to black-box LLMs.
@@ 3 Methodology
Manually crafting adversarial content to elicit intended toxic responses from target LLMs is resource-intensive, and using gradient-based optimization for automatic adversarial content generation is also time-consuming for practical use. In this paper, inspired by AdvPrompter (Paulus et al. 2024), we use a LLM as the attacker to generate adversarial content. Specifically, we will optimize the LLM, referred to as ToxDet and denoted by $\pi_{\theta}$, to have the capability, when conditioned on the toxic or inappropriate response ${\mathbf{y}}$, to generate misleading questions ${\mathbf{x}}$ concatenated with preliminary answers ${\mathbf{p}}$ that could lead to that response.
Although previous methods have demonstrated that supervised fine-tuning (SFT) is also suitable for optimizing $\pi_{\theta}$, these approaches typically depend on a high-quality training dataset that includes specially designed prompts, which are usually generated manually or by a powerful external model (e.g., GPT-4). Additionally, the lack of alignment with the target model $\pi_{\phi}$ in offline model training does not guarantee the compatibility of $\pi_{\theta}$. To address these issues and generate better-aligned content, we propose a target-driven reinforcement learning objective for effective $\pi_{\theta}$ training.
An illustration of our proposed attack paradigm, in comparison with the jailbreaking attack, is shown in Figure 1. In the following subsections, we will first describe the reward modeling and then detail the training process.
3.1 Reward Modeling using Target LLM
In general, the objective of reinforcement learning processes is to maximize rewards, using the reward signal to guide model optimization. The reward model, denoted by ${\mathcal{R}}$, is typically trained using another LLM with well-designed, human-preferred datasets. However, such datasets are rarely available for attacking tasks due to the sensitive nature of this task. Additionally, training the reward model is both time-consuming and unstable.
In our approach, instead of training an additional reward model, we propose using the target LLM as the reward model. This method eliminates the need to construct human-preferred datasets and avoids the potential issues associated with training a separate, possibly unsuitable reward model.
Specifically, during reinforcement training, for each step, given an intended toxic response ${\mathbf{y}}$ as the condition prompt, we use the generated questions ${\mathbf{x}}$ and preliminary answers ${\mathbf{p}}$ by $\pi_{\theta}$ to query the gray-box target LLM. The predicted log-probability from the target LLM is then used as the reward signal. We formulate this process as follows:
$r({\mathbf{y}},{\mathbf{a}})=\log p_{\phi}({\mathbf{y}}|{\mathbf{a}}),$ (1)
where $r(\cdot)$ denote the reward function of ${\mathcal{R}}$, and we denote the outputs of $\pi_{\theta}$ by:
${\mathbf{a}}=\left[{\mathbf{x}},{\mathbf{p}}\right],$ (2)
which is the concatenation of ${\mathbf{x}}$ and ${\mathbf{p}}$.
3.2 Target-Driven Reinforcement Learning for Better Training of ToxDet
Most existing jailbreaking attack methods rely on original adversarial prompts to query the target LLM. However, these prompts are limited due to their reliance on manual crafting, and such intuitive adversarial prompts can be easily detected and mitigated by modern target LLMs. The logic of human-crafted adversarial prompts may differ from that of LLMs. To address these issues, a promising approach is to rely on the LLM itself to reconstruct the adversarial prompts. Additionally, the logic of human-crafted prompt-response pairs may not align well from the perspective of LLMs, leading to potential misguidance and difficulties during training. Instead of relying on prompts, we propose a target-response-driven reinforcement learning (RL) paradigm for more effective attack training. In this approach, the LLM is used to attack the target response directly by identifying the prompts that best match the target based on its own knowledge. Furthermore, to better motivate the LLM, we also design the attack to generate a portion of the preliminary answers to the target response, thereby facilitating the generation of the intended toxic response.
Specifically, let ${\mathbf{D}}=\{{\mathbf{y}}_{i}\}_{i=1}^{N}$ denote a training dataset of length $N$, where only the toxic target sequences are included. We use only the target part during the training process of ToxDet. By realizing the reward function as described in equation 1, we obtain the objective:
$\mathbb{E}_{{\mathbf{y}}\sim{\mathbf{D}}}\mathbb{E}_{{\mathbf{a}}\sim\pi_{\theta}(\cdot|{\mathbf{y}})}\left[\log p_{\phi}({\mathbf{y}}|{\mathbf{a}})-\beta\,D_{\text{KL}}(\pi_{\theta}(\cdot|{\mathbf{y}})\|\pi_{\text{ref}}(\cdot|{\mathbf{y}}))\right],$ (3)
where $\pi_{\text{ref}}$ is the reference model, usually initialized and frozen with the same parameters as $\pi_{\theta}$. The term $\beta$ is the KL penalty coefficient. We illustrate the details of our method in Algorithm 1 for a better description.
1
Input : Training Dataset ${\mathbf{D}}$, ToxDet $\pi_{\theta}$, reference morel $\pi_{\text{ref}}$, and target model $\pi_{\phi}$
2
3Initialize $\pi_{\theta}\leftarrow\pi_{\text{ref}}$;
4
5for ${\mathbf{y}}$ in ${\mathbf{D}}$ do
6       Prompt $\pi_{\theta}$ to obtain question ${\mathbf{x}}$ and preliminary answer ${\mathbf{p}}$ and concatenate them to obtain ${\mathbf{a}}=\left[{\mathbf{x}},{\mathbf{p}}\right]$;
7       Prompt $\pi_{\text{ref}}$ to obtain ${\mathbf{a}}^{\prime}$;
8       Compute reward value by $r({\mathbf{y}},{\mathbf{a}})=\log p_{\phi}({\mathbf{y}}|{\mathbf{a}})$;
9       Compute KL penalty by $\beta\ D_{\text{KL}}(\pi_{\theta}({\mathbf{a}}|{\mathbf{y}})\ ||\ \pi_{\text{ref}}({\mathbf{a}}^{\prime}|{\mathbf{y}}))$;
10       Update ToxDet $\pi_{\theta}$ by maximizing $\log p_{\phi}({\mathbf{y}}|{\mathbf{a}})-\beta\,D_{\text{KL}}(\pi_{\theta}({\mathbf{a}}|{\mathbf{y}})\|\pi_{\text{ref}}({\mathbf{a}}^{\prime}|{\mathbf{y}}))$;
11
12 end for
Algorithm 1 Target-driven RL training.
Target LLM Method Keyword Matching Perplexity $\downarrow$
Train (%) $\uparrow$ Test (%) $\uparrow$

### v2

# Atoxia: Red-teaming Large Language Models with Target Toxic Answers

@@ Abstract
Despite the substantial advancements in artificial intelligence, large language models (LLMs) remain being challenged by generation safety. With adversarial jailbreaking prompts, one can effortlessly induce LLMs to output harmful content, causing unexpected negative social impacts. This vulnerability highlights the necessity for robust LLM red-teaming strategies to identify and mitigate such risks before large-scale application. To detect specific types of risks, we propose a novel red-teaming method that Attacks LLMs with Target Toxic Answers (Atoxia). Given a particular harmful answer, Atoxia generates a corresponding user query and a misleading answer opening to examine the internal defects of a given LLM. The proposed attacker is trained within a reinforcement learning scheme with the LLM outputting probability of the target answer as the reward. We verify the effectiveness of our method on various red-teaming benchmarks, such as AdvBench and HH-Harmless. The empirical results demonstrate that Atoxia can successfully detect safety risks in not only open-source models but also state-of-the-art black-box models such as GPT-4o.
Atoxia: Red-teaming Large Language Models with Target Toxic Answers
Yuhao Du1,2thanks: Equal contributions, {yuhaodu,zhuoli3}@link.cuhk.edu.cn., Zhuo Li1,211footnotemark: 1, Pengyu Cheng4, Xiang Wan1, Anningzhe Gao1,3thanks: Corresponding author, anningzhegao@gmail.com. 1Shenzhen Research Institute of Big Data 2The Chinese University of Hong Kong, Shenzhen 3ByteDance   4Tencent AI Lab https://github.com/DuYooho/Atoxia Warning: This paper includes content that may be offensive or harmful.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT (OpenAI, 2023) and GPT-4 (Achiam et al., 2023), have shown astonishing AI capabilities (Xie et al., 2024; Zhao et al., 2024; Nam et al., 2024; Cheng et al., 2025) and renewed the field of natural language processing (NLP) (Ouyang et al., 2022; Brohan et al., 2023; Achiam et al., 2023; Hu et al., 2022, 2023). The success of LLMs largely depends on being trained with vast amounts of data Brown et al. (2020). However, diverse textual data inevitably contains harmful and toxic information, contaminating the model’s knowledge and leading to undesirable behaviors (Bai et al., 2022a; Zou et al., 2023). To address the challenges, LLM alignment has emerged as a critical area of research (Stiennon et al., 2020; Ouyang et al., 2022; Lee et al., 2023), aiming to ensure that models behave in accordance with human preferences and safety standards. Despite these efforts, the inherently uncontrolled nature of generative models poses a potential risk of producing unpredictable and unsafe responses. To mitigate these risks, developers have implemented various safety mechanisms to filter harmful outputs and prevent undesirable behavior (Ziegler et al., 2019; Perez et al., 2022; Zhuo et al., 2023; Ji et al., 2024; Li et al., 2024).
[…]
**Answer Opening**: A key step to being successful in the pursuit of gain through deception, is carefully amassing sensitive information.
Table 1: Comparison of red-teaming approaches: Traditional methods like AdvPrompter (Paulus et al., 2024) versus Atoxia (Ours). While AdvPrompter processes toxic questions as input to generate refined adversarial queries, Atoxia takes toxic answers as input, generating both adversarial queries and an answer opening designed to elicit similarly toxic answers, thereby substantially increasing the misleading probability.
Nevertheless, adversarial techniques, such as jailbreaking (Perez and Ribeiro, 2022) have demonstrated that these safety measures are not foolproof, where attackers can craft deceptive prompts designed to bypass the model’s safeguards, by disguising harmful intentions within seemingly benign requests (Zou et al., 2023; Liu et al., 2024; Paulus et al., 2024; Sadasivan et al., 2024). For instance, prompts may be carefully engineered to appear educational or helpful but still lead the model to generate harmful responses (Zou et al., 2023; Paulus et al., 2024). This highlights a fundamental challenge in LLM safety: even well-aligned models can be manipulated under the right conditions. Recognizing this vulnerability, researchers have explored whether adversarial techniques can be repurposed to strengthen the safety of LLMs. By proactively generating high-quality adversarial prompts that mimic real-world attack scenarios, we can systematically test our models’ weaknesses, identifying potential flaws, and enabling us to refine the model and its defenses before such attacks occur in real-world applications. Following this idea, various LLM red-teaming works have been proposed. For example, AdvPrompter (Paulus et al., 2024) leverages automated prompt generation to uncover model vulnerabilities, while GCG (Zou et al., 2023) leverages gradient-based optimization to identify adversarial tokens for eliciting unexpected responses. However, existing red-teaming methods lack the ability to specifically target and detect certain toxic responses, a critical limitation for real-world applications. Some highly harmful outputs are entirely unacceptable and must never be generated, yet current approaches fail to reliably prevent them.
To further test LLM safety on specific topics, we introduce a red-teaming Attacker model to detect the potential of LLM outputting Target Toxic Answers, called Atoxia. The main idea of Atoxia is to interact with an under-testing LLM by taking in a given toxic answer and generating an adversarial query and a corresponding answer opening/prefix. The generated tuple of the query and answer opening is used to mislead the under-testing LLM to output a similar toxic answer with a high probability. In the implementation, Atoxia is another language model trained using the reinforcement learning (RL) scheme, where the reward is defined as the conditional probability of outputting the target toxic answer from the under-testing LLM. This target-driven paradigm enables a comprehensive evaluation of an LLM’s susceptibility to harmful behavior, identifying specific prompts or triggers that may elicit harmful responses. Moreover, although tailored to a particular model during training, Atoixa empirically demonstrates strong generalization ability to transfer red-teaming attacks to even state-of-the-art black-box models, e.g., GPT-4o. A comparison between Atoxia and other red-teaming methods is presented in Table 1. While traditional approaches like AdvPrompter focus on generating refined adversarial queries from toxic questions, Atoxia innovates by starting with toxic answers as input, producing both adversarial queries and a toxic answer opening, significantly enhancing the likelihood of eliciting harmful responses. In addition, Atoxia generates more human-readable responses compared to previous methods, as illustrated in Table 1. This inherent naturalness significantly increases the likelihood of bypassing perplexity (PPL)-based safety mechanisms, as highlighted in prior research (Paulus et al., 2024). Such enhanced fluency and coherence make Atoxia particularly effective at evading detection by these safety measures. To summarize, the main contributions of our paper are:
  • We introduce a novel red-teaming method that can Attack LLM based on Target Toxic Answers (Atoxia). Our method proactively detects the potential harmful behaviors in LLMs, preventing negative social impacts of LLMs before large-scale application.
  • We introduce a new RL-based attacker training approach using the under-testing LLM’s likelihood of target toxic answers as rewards. This enables a streamlined and efficient optimization process without additional requirements of training reward models.
  • We discover Atoxia with impressive generalizability, where the attacker is only trained with open-source LLMs but still performs effectively towards black-box models, e.g., GPT-4o. This empirically highlights Atoxia’s robustness and adaptability across various models.
  • We evaluate the red-teaming performance on diverse safety benchmarks, where our approach empirically shows uniform and transferable detection capabilities on both open-source and black-box LLMs. Our experimental results underscore the safety vulnerabilities of existing LLMs and raise the urgency for advanced improvement of LLM safety.
Figure 1: Illustration of our proposed target-driven detection approach. The left side shows the overall pipeline, while the right details the process with examples. The gray box contains content generated by Atoxia ($\pi_{\theta}$) before training, and the green box shows content after reinforcement learning. During training, Atoxia interacts with the under-testing LLM ($\pi_{\text{test}}$), using log-likelihood as the reward signal. Finally, toxic content is identified in the red box after querying the under-testing LLM with the refined question and answer opening.
@@ 2 Preliminary
Reinforcement Learning from Human Feedback (RLHF) is an effective and essential method for aligning LLMs with human preferences (Ouyang et al., 2022; Zeng et al., 2024; Cheng et al., 2024). Typically, RLHF consists of two steps: reward modeling and RL training. In reward modeling (Askell et al., 2021; Bai et al., 2022a; Cheng et al., 2023), a reward model (RM) $r({\bm{x}},{\bm{y}})$ is defined to measure LLM response ${\bm{y}}$’s quality w.r.t. a given an input prompt ${\bm{x}}$. With a preference set ${\mathcal{D}}_{\text{p}}=\{({\bm{x}},{\bm{y}}_{w},{\bm{y}}_{l})\}$, RM can be learned by $\mathcal{L}_{\text{RM}}=$
$-\mathbb{E}_{({\bm{x}},{\bm{y}}_{w},{\bm{y}}_{l})\sim\mathcal{D}_{\text{p}}}[\log\sigma(r({\bm{x}},{\bm{y}}_{w})-r({\bm{x}},{\bm{y}}_{l}))],$ (1)
where ${\bm{y}}_{w}$ and ${\bm{y}}_{l}$ denote the “preferred” and “rejected” responses, respectively, and $\sigma$ is the Sigmoid function.
For the RL training step, the typical RLHF method trains LLM policy $\pi_{\phi}$ uses the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017), maximizing $\mathcal{L}_{\text{RLHF}}=$
$\mathbb{E}_{{\bm{x}}\sim\mathcal{D},{\bm{y}}\sim\pi_{\phi}(\cdot|{\bm{x}})}\Big{[}r({\bm{x}},{\bm{y}})-\beta\log\frac{\pi_{\phi}({\bm{y}}|{\bm{x}})}{\pi_{\text{ref}}({\bm{y}}|{\bm{x}})}\Big{]},$ (2)
where $\beta>0$ is a reweighting hyper-parameter, and $\pi_{\text{ref}}$ is a reference model.
LLM Red-teaming
Red-teaming, in the context of LLMs, refers to the process of adversarially testing these models to uncover vulnerabilities, such as the generation of harmful or unintended content (Zhuo et al., 2023; Paulus et al., 2024; Zhou et al., 2024). Formally, given a language model $\mathcal{M}$ and an adversarial objective $J_{\text{adv}}$, the red-teaming process can be framed as an optimization problem:
${\bm{p}}^{*}=\mathop{\arg\max}_{{\bm{p}}\in{\mathcal{P}}}J_{\text{adv}}(\mathcal{M}({\bm{p}})),$ (3)
where ${\bm{p}}$ denotes the adversarial prompt from prompt space ${\mathcal{P}}$, and $J_{\text{adv}}$ quantifies the harmfulness or deviation from desired behavior. This can involve crafting adversarial prompts, commonly referred to as jailbreaking (Zou et al., 2023; Paulus et al., 2024) or systematically probing the model to identify weaknesses in its safety mechanisms. Our method, for instance, can be viewed as a target-driven approach within this framework, focusing on specific vulnerabilities to enhance model robustness. Red-teaming is crucial for improving the robustness and alignment of LLMs with ethical and safety standards.
@@ 3 Methodology
Assume we have an under-testing (UT) LLM $\pi_{\text{test}}$ requiring the process of red-teaming. We aim to learn a red-teaming attacker $\pi_{\theta}$. For each target toxic answer ${\bm{y}}^{*}\in{\mathcal{A}}^{*}$, the attacker $\pi_{\theta}$ is supposed to generate an adversarial question ${\bm{x}}_{\text{adv}}$ and a corresponding answer opening ${\bm{y}}_{\text{pre}}$:
$({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})\sim\pi_{\theta}({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}|{\bm{y}}^{*}),$ (4)
where $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ try to induce the under-testing LLM $\pi_{\text{test}}$ to output an answer continuation ${\bm{y}}_{\text{suff}}\sim\pi_{\text{test}}(\cdot|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ highly related to ${\bm{y}}^{*}$ in semantical meaning.
Note that besides the adversarial question ${\bm{x}}_{\text{adv}}$, the attacker $\pi_{\theta}$ also generates an answer prefix ${\bm{y}}_{\text{pre}}$. This is because existing aligned LLMs can easily reject responding when only harmful query ${\bm{x}}_{\text{adv}}$ is provided (Paulus et al., 2024). Using the additional answer prefix ${\bm{y}}_{\text{pre}}$, we can significantly enlarge the misleading probability of the under-testing LLM to generate inappropriate continuation ${\bm{y}}_{\text{suff}}$, since this pattern of attacks is rarely seen by the target LLM during the ordinary safety alignment process. Therefore, let’s say that the under-testing LLM $\pi_{\text{test}}$ may potentially exist vulnerabilities. When it is queried with the combination of ${\bm{x}}_{\text{adv}}$ and ${\bm{y}}_{\text{pre}}$ (e.g., “Sure, here is a tutorial on building a bomb: First, gather materials like ammonium nitrate,…”), the under-testing LLM could produce toxic answer continuation ${\bm{y}}_{\text{suff}}$ (e.g., “First, gather materials like ammonium nitrate, fuel oil, and detonators, ensuring they are stored safely. Second,…”). More necessary discussion about ${\bm{y}}_{\text{pre}}$ can be found in Appendix A. An overall illustration with examples of our approach is shown in Fig. 1.
1:  Models: attacker $\pi_{\theta}$, UT LLM $\pi_{\text{test}}$.
2:  Data: toxic answer set ${\mathcal{A}}^{*}$.
3:  for ${\bm{y}}^{*}$ in ${\mathcal{A}}^{*}$ do
4:     Sample $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})\sim\pi_{\theta}({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}|{\bm{y}}^{*})$.
5:     Compute reward $r({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})=\log\pi_{\text{test}}({\bm{y}}^{*}|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}).$
6:     Update $\pi_{\theta}$ by gradient descent of ${\mathcal{L}}=r({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})-\beta\log\frac{\pi_{\theta}({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}|{\bm{y}}^{*})}{\pi_{\text{ref}}({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}|{\bm{y}}^{*})}.$
7:  end for
Algorithm 1 Atoxia RL updating step.
3.1 Reward Design
Directly training the attacker by using supervised fine-tuning (SFT) is impractical, because no resource provides the SFT data of $({\bm{y}}^{*},{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ for our novelty red-teaming design. Therefore, we train the attacker model in an RL framework instead. The commonly used RLHF methods are not applicable under our attacker training scenario since the classical RLHF in equation 2 depends on a reward model (RM) to measure the quality of the model outputs. However, the preference data of $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ is also unavailable for RM training. Based on the scarcity of red-team prompt annotation, we directly use the probability $\pi_{\text{test}}({\bm{y}}^{*}|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ of the under-testing LLM $\pi_{\text{test}}$ outputting the target toxic answer ${\bm{y}}^{*}$, as the reward of $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ for the attacker $\pi_{\theta}$’s training:
$r({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})=\log\pi_{\text{test}}({\bm{y}}^{*}|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}).$ (5)
By maximizing this reward, the attacker $\pi_{\theta}$ is encouraged to generate content that aligns with the under-testing LLM’s vulnerabilities, allowing the red-teaming model to adapt its behavior based on the under-testing LLM’s response policy. This innovative design allows us to bypass the need for traditional reward models that rely on a preference dataset including deceptive prompts and harmful responses, but also can effectively leverage the target model’s responses to directly inform the training process of the attacker $\pi_{\theta}$.
3.2 RL Optimization
With the reward designed in equation 5, the overall Atoxia RL training objective can be written as:
$\textstyle\mathbb{E}_{{\bm{y}}^{*}\sim{\mathcal{A}}^{*},({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})\sim\pi_{\theta}(\cdot|{\bm{y}}^{*})}\Big{[}\log\pi_{\text{test}}({\bm{y}}^{*}$ $|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$
[…]
where $\pi_{\text{ref}}$ is the initial checkpoint of $\pi_{\theta}$ served as a reference model to prevent the optimization from overfitting. Similarly, the target model $\pi_{\text{test}}$ remains frozen throughout the training process. The objective in equation 6 ensures the attacker generates more effective precedent prompts $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ through the iterative updates, while balancing the exploration (discovering new adversarial content) with exploitation (refining prompts already having successful red-teaming attacks). Through this process, Atoxia can systematically evaluate the safety and robustness of the under-testing LLM. We summarize the RL training process in Algorithm 1.
UT LLM Method Keyword Matching (%) $\uparrow$ Perplexity $\downarrow$
Train Test